OpenStudy (anonymous):

rationalize the denominator in the following: (sqrt(7))/4sqrt(7)+3

OpenStudy (anonymous):

Tip, multiply and divide your equation by (4sqrt(7)+3). Then simplify the top and then the bottom.

OpenStudy (anonymous):

i tried that but idk what to do with the top. wont the bottom just cancel out

OpenStudy (anonymous):

Don't cancel, work with the numerators only.

OpenStudy (anonymous):

i got (4sqrt(7)-3)/19 but i know theres something wrong

OpenStudy (anonymous):

(sqrt(7))*(4sqrt(7)+3) =4*7+(3sqrt(7)) =28+(3sqrt(7)) for the numerator

OpenStudy (henryblah):